Please can you help me with the equation to solve the following word problem:
A Chemist is making a new cleaning product by mixing 2 solutions. How many litres of a 35% solution should be added to the 15% solution to obtain 8 litres of a cleaning solution with 20% concentration?
Thanking you in advance for your assistance.

Answer by Earlsdon(6291)

Let x = the required number litres of the 35% solution then (8-x) will be the required number of litres of the 15% solution.
Note that x + (8-x) = 8, so, after converting the percentages to their decimal equivalents, we have...
0.35x+0.15(8-x) = 0.2(8) Simplify and solve for x.
0.35x+1.2-0.15x = 1.6
0.2x = 0.4
x = 2.
The chemist will need to add 2 litres of the 35% solution to (8-2=6) 6 litres of the 15% solution to obtain 8 litres of 20% solution.